Circular No. 8083

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002kl Further to IAUC 8081, T. Dahlen reports the discovery of another

apparent supernova on ACS images of the Hubble Deep Field North, located at α = 12^{h}37^{m}49^{s}.30, δ = +62^{o}14′06′′.1, which is 0′′.15 east and 0′′.45 south of the center of the nearest visible galaxy. Available z-band magnitudes for SN 2002kl: 2002 Nov. 21.76, 24.8; 2003 Jan. 2.91 UT, 25.0; Feb. 21.95, [28.

SUPERNOVAE 2003V, 2003ah, 2003av, 2003ax, 2003bf A. V. Filippenko and R. Chornock, University of California, Berkeley,

report that inspection of CCD spectra (range 310–1000 nm) obtained on Feb. 27 UT with the Keck I 10-m telescope (+ LRIS) shows that SN 2003V (IAUC 8060) is of type Ia, ∼ 2 months past maximum brightness. SN 2003av (IAUC 8077) and SN 2003ax (IAUC 8079) are also of type Ia, ∼ 2 weeks past maximum. Yet another type-Ia supernova is SN 2003ah (IAUC 8068), ∼ 12 days past maximum. SN 2003bf (cf. IAUC 8082) is of type Ia, very close to maximum brightness.

COMET C/2002 Y1 (JUELS-HOLVORCEM) D. K. Lynch, R. W. Russell, and D. L. Kim, The Aerospace Corpora-

tion; M. L. Sitko, University of Cincinnati; and R. B. Perry, Langley Research Center, NASA, report that 3–14-μm spectroscopy of comet C/2002 Y1, obtained on Feb. 20.6 UT with the NASA Infrared Telescope Facility 3-m telescope (+ Aerospace Broadband Array Spectrograph System), yielded a narrowband magnitude of N [10.2 μm] = 3.5 ± 0.1. The spectrum shows a smooth featureless continuum with an 8–13-μm color temperature of ∼ 280 ± 20 K, roughly 12 times higher than the radiative equilibrium blackbody temperature. The comet was not detected between 3 and 8 μm, and an upper limit to the silicate emission feature was ≈ 10 percent of the 8–13-μm continuum.

YY HERCULIS E. O. Waagen, AAVSO, reports that this variable is in outburst for the

first time since May 1993 (cf. IAUC 5805; it peaked at m_{v} = 11.3 in July 1993, slowly faded over a year to m_{v} = 13.6, rebrightened to m_{v} = 12.3 in early 1995, and then declined to its normal minimum near mag 13 by Sept. 1995). Recent CCD V magnitudes from R. James, Las Cruces, NM: Feb. 2.518 UT, 12.9; 9.513, 12.8; 17.496, 12.1; 23.503, 11.9; 27.472, 11.9.

2003 February 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green